An insurance company claims that for x thousand​ policies, its monthly revenue in dollars is given by R(x)=250x and its monthly cost in dollars is given by C(x)=200x+20,000.

a. Find the​ break-even point.

b. Graph the revenue and cost equations on the same axes.

c. From the​ graph, estimate the revenue and cost when

x=200.

a. The​ break-even point is = .

b. Graph the revenue and cost equations on the same axes.

c. From the​ graph, estimate the revenue and cost when

x=200. the revenue is ​ $ When x=200 the cost is ​ $    .
Answer #1

Solna given A kla) = 2500 () 2008 + 20000 @ for break even point RIN) = c(0) 250x = doont 20000 3 503 = V0000 1 x=400 © RIX)| Revenue when (n=200) = $50,000 <(x) = 2000 + POOVU Put a= 200 (14) = 200 x 200 + 20000 (LX) = 40000 + 20000 ((x) = 60000 Co